# RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis

### v1

@@ V Conclusion and Future Work
We introduce RobotFingerPrint (RFP), a novel and effective way to unify grasp representations across different grippers into a common coordinate space. The unified grasp representation is dense in nature, requires no manual specification and has an unambiguous mapping to each gripper’s individual fingers. Such a mapping helps in synthesizing diverse and stable grasping configurations via a hand-object correspondence and well-initialized grasp optimization which results in strong performance across several evaluation scenarios. We plan to extend the core idea of RFP-driven grasp synthesis to partially observed and incomplete point clouds in real-world grasping scenarios and also utilizing it for grasp transfer from humans. Another avenue of exploration would be to investigate a learned reverse mapping from a coordinate map to grasp prediction without the intermediate grasp optimization.
ACKNOWLEDGMENT
@@
This work was supported by the Sony Research Award Program and the National Science Foundation (NSF) under Grant No. 2346528.

### v2

@@ VII Conclusion and Future Work
We introduce RobotFingerPrint (RFP), a novel and effective way to unify grasp representations across different grippers into a common coordinate space. The unified grasp representation requires no manual specification and has a precise, dense mapping for each gripper’s individual fingers. This helps in synthesizing diverse and stable grasp configurations via (1) hand-object correspondence and (2) well-initialized grasp pose. RFP shows strong performance across several evaluation scenarios in both simulation and real-world testing with unseen Fetch gripper. Further, it enables direct grasp transfer between novel grippers without manual labeling and enables downstream applications in learning from human demonstrations. Future work would be to address a potential limitation in collision-aware grasp generation on partial object observations, and real-time grasp inference for use in a closed-loop setting.
@@
Acknowledgment. This work was supported by the Sony Research Award Program and the National Science Foundation (NSF) under Grant No. 2346528.